Solve the given system by the substitution method.
-9x+y=43
6x+18y=102

Answer by waynest(281)

-9x + y = 43
6x + 18y = 102
y = 9x + 43
substitute 9x + 43 for y in the second equation
6x + 18(9x + 43) = 102
6x + 162x + 774 = 102
168x + 774 - 774 = 102 - 774
168x = -672
168x/168 = -672/168
x = -4
substitute -4 for x in equation 1
-9(-4) + y = 43
36 + y = 43
36 - 36 + y = 43 - 36
y = 7
check with both equations:
-9(-4) + 7 = 43
36 + 7 = 43
43 = 43
6(-4) + 18(7) = 102
-24 + 126 = 102
102 = 102